### v1

# KYUSHU-HET-290

Abstract
Employing the modified Villain lattice formulation of the axion QED, we present an alternative and much simpler derivation of the conclusion of Ref. Honda:2024sdz that the sweep of the axial $U(1)$ non-invertible symmetry operator over the (non-genuine) gauge invariant ’t Hooft line operator with an integer magnetic charge does not leave any effect. The point is that such a ’t Hooft line can be represented by a boundary of a (non-topological) defect that is invariant under the axial transformation on the axion field.
[…]
B01,B04,B31
1 Introduction and summary
The study in Refs. Choi:2022jqy ; Cordova:2022ieu pointed out the possibility that the axial $U(1)$ rotation of the fermion field with fractional angles, although it suffers from the ordinary axial $U(1)$ anomaly, may be understood as a generalized form of the symmetry Gaiotto:2014kfa (see Refs. Schafer-Nameki:2023jdn ; Bhardwaj:2023kri ; Shao:2023gho for reviews), the non-invertible symmetry Aasen:2016dop ; Bhardwaj:2017xup ; Chang:2018iay ; Thorngren:2019iar ; Komargodski:2020mxz ; Koide:2021zxj ; Choi:2021kmx ; Kaidi:2021xfk ; Hayashi:2022fkw ; Choi:2022zal ; Kaidi:2022uux ; Roumpedakis:2022aik ; Bhardwaj:2022yxj ; Cordova:2022ieu ; Choi:2022jqy ; Bhardwaj:2022lsg ; Karasik:2022kkq ; GarciaEtxebarria:2022jky ; Choi:2022fgx ; Yokokura:2022alv ; Nagoya:2023zky ; Anber:2023mlc . In Ref. Honda:2024sdz , the present authors studied how the symmetry operator of the axial $U(1)$ non-invertible symmetry acts on the (non-genuine) ’t Hooft line operator by modeling the axial anomaly by the axion and employing the modified Villain lattice formulation of $U(1)$ gauge theory Sulejmanpasic:2019ytl (see also Ref. Gorantla:2021svj ), which allows a straightforward introduction of the dual $U(1)$ gauge field. In Ref. Honda:2024sdz , it was concluded that the sweep of the symmetry operator over the ’t Hooft line operator does not leave any effect. This conclusion appears inequivalent with the phenomenon concluded in Refs. Choi:2022jqy ; Cordova:2022ieu that the sweep leaves non-trivial traces. Although we believe that under assumptions made in Ref. Honda:2024sdz , our conclusion is correct, the argument in Ref. Honda:2024sdz might be somewhat complicated and, in particular, it appears that the conclusion depends crucially on a miraculous cancellation between the effect of the dressing factor in the ’t Hooft line and the would-be Witten effect Witten:1979ey arising from the axion topological coupling.
In this paper, we present an alternative and much simpler derivation of the conclusion in Ref. Honda:2024sdz , although the basic assumptions are common. Considering the basic importance of this issue in possible applications of the axial $U(1)$ non-invertible symmetry, especially in the form of the selection rules, we believe that it is useful to do this. In this paper, we show that the conclusion in Ref. Honda:2024sdz can be derived from the following observation: The (non-genuine) ’t Hooft line operator considered in Ref. Honda:2024sdz can be realized by a background 2-cocycle gauge field, which couples to the (non-conserved) electric 2-cocycle current $\star f$. Since this coupling is not influenced by the axial transformation on the axion field, $\phi\to\phi+\alpha$, the symmetry operator does not cause any effect on the ’t Hooft line. This simple observation explains the above miraculous cancellation. Also, we hope that the present simple argument gives an idea that our conclusion, although it is shown by employing a particular lattice action, is quite insensitive on the details of the lattice action one adopts. The extension of the present argument to the case of the action of the 1-form non-invertible symmetry in the axion QED Choi:2022fgx ; Yokokura:2022alv , the action of the 1-form non-invertible symmetry operator on the axion string, would be an interesting problem. [^1] Since the (non-genuine) gauge invariant axion string can be represented by a background 1-cochain gauge field introduced in the kinetic term of the axion as we will see in Sect. 2.5, and the 1-form non-invertible symmetry is associated with the shift of the gauge potential, repeating a similar argument as that in the present paper, we expect that the sweep of the symmetry operator over the axion string does not leave any effect.
2 The action of the axial $U(1)$ non-invertible symmetry operator
2.1 Lattice action
We take the lattice action of Ref. Honda:2024sdz (Eq. (2.24) there) as an explicit starting point for our argument. It is based on the modified Villain formulation of the $U(1)$ lattice gauge theory Sulejmanpasic:2019ytl . After the integration by parts on the periodic hypercubic lattice $\Gamma$, the lattice action can be written as [^2] We denote the hypercube, the cube, the plaquette, and the link on $\Gamma$ by $h$, $c$, $p$, and $l$, respectively. The $p$-cochain, the cup product $\cup$, and the coboundary operator $\delta$ on $\Gamma$ are, roughly speaking, lattice analogues of the $p$-form, the wedge product $\wedge$, and the exterior derivative $\mathrm{d}$ on $T^{4}$, respectively. For precise definitions of the $p$-cochain, the cup product, the higher cup product $\cup_{1}$, the coboundary operator and the “Hodge dual” $\star$ on the hypercubic lattice, see Refs. Chen:2021ppt ; Jacobson:2023cmr . On the cup product, the coboundary operator $\delta$ satisfies the Leibniz rule, $\delta(\alpha\cup\beta)=\delta\alpha\cup\beta+(-1)^{p}\alpha\cup\delta\beta$, where $\alpha$ is a $p$-cochain. The cup product is not commutative on the cochain level but the commutator is expressed by the higher cup as $(-1)^{pq}\beta\cup\alpha=\alpha\cup\beta+(-1)^{p+q}[\delta(\alpha\cup_{1}\beta)-\delta\alpha\cup_{1}\beta-(-1)^{p}\alpha\cup_{1}\delta\beta]$, where $\beta$ is a $q$-cochain.
$S=\sum_{h\in\Gamma}\left\{\frac{1}{g_{0}^{2}}f\cup\star f+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi-i\frac{\mathrm{e}^{2}}{8\pi^{2}}\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]\right\}.$ (2.1)
In this expression,
$f:=\delta a+2\pi z,$ (2.2)
is the $U(1)$ field strength 2-cochain ($a\in\mathbb{R}$ is the $U(1)$ gauge potential 1-cochain, $\delta$ is the coboundary operator and $z\in\mathbb{Z}$ is an integer 2-cochain field). $\tilde{a}\in\mathbb{R}$ is a 1-cochain Lagrange multiplier field and the integration over $\tilde{a}$ imposes the Bianchi identity $\delta z=0$; $\tilde{a}$ can be regarded as the dual $U(1)$ gauge potential. For the axion scalar field $\phi$, $\partial\phi$ is the combination defined by
$\partial\phi:=\delta\phi+2\pi\ell,$ (2.3)
where $\phi\in\mathbb{R}$ is a 0-cochain and $\ell\in\mathbb{Z}$ is an integer 1-cochain field. $\chi\in\mathbb{R}$ in Eq. (2.1) is another 2-cochain auxiliary field which imposes the another Bianchi identity $\delta\ell=0$. The lattice $U(1)$ Chern–Simons form $(\mathrm{CS})$ is defined by Jacobson:2023cmr  [^3] The motivation for this definition may be understood as follows: In the present lattice formulation, $f\cup f$ is not a total divergence. One can make this total divergence up to $8\pi^{2}\mathbb{Z}$ by adding terms containing $\delta z$ as $f\cup f-4\pi a\cup\delta z+2\pi f\cup_{1}\delta z=\delta(\mathrm{CS})+4\pi^{2}P_{2}(z).$ (2.4) That is, under the Bianchi identity $\delta z=0$, the $\theta$ term defined by $(\theta/8\pi^{2})f\cup f=(\theta/8\pi^{2})[\delta(\mathrm{CS})+4\pi^{2}P_{2}(z)]$ is manifestly topological and changes by $2\pi\mathbb{Z}$ under $\theta\to\theta+2\pi$.
$(\mathrm{CS}):=a\cup f+2\pi z\cup a+2\pi a\cup_{1}\delta z.$ (2.5)
Finally, the Pontryagin square is defined by
$P_{2}(z):=z\cup z+z\cup_{1}\delta z.$ (2.6)
In the modified Villain formulation Sulejmanpasic:2019ytl , the lattice action and observables must be invariant under the following three gauge transformations: The $\mathbb{Z}^{(1)}$ gauge transformation,
$a\to a+2\pi m,\qquad z\to z-\delta m,$ (2.7)
where $m\in\mathbb{Z}$ is a 1-cochain. The $\mathbb{Z}^{(0)}$ gauge transformation,
$\phi\to\phi+2\pi k,\qquad\ell\to\ell-\delta k,$ (2.8)
where $k\in\mathbb{Z}$ is a 0-cochain. The $\mathbb{R}^{(0)}$ gauge transformation,
$a\to a+\delta\lambda,\qquad z\to z,$ (2.9)
where $\lambda\in\mathbb{R}$ is a 0-cochain. The first two gauge invariances remove redundancies in the above description of the lattice $U(1)$ gauge theory with a $2\pi$ periodic scalar field. The last one is the ordinary local $U(1)$ gauge symmetry.
In the above form of the lattice action (2.1), the invariance under the $\mathbb{Z}^{(0)}$ gauge transformation (2.8) is manifest, because $\partial\phi$ in Eq. (2.3) is a manifestly $\mathbb{Z}^{(0)}$ gauge invariant combination. It can be seen that Honda:2024sdz the lattice action (2.1) is invariant also under other two gauge transformations, [^4] One finds that, under the $\mathbb{Z}^{(1)}$ gauge transformation (2.7), $(\mathrm{CS})$ $\to(\mathrm{CS})-2\pi\delta(m\cup a)+8\pi^{2}m\cup z-4\pi^{2}\delta(m\cup_{1}z)+4\pi^{2}\delta m\cup_{1}z-4\pi^{2}\delta m\cup m,$ $P_{2}(z)$ $\to P_{2}(z)-2\delta m\cup z-\delta(\delta m\cup_{1}z)+\delta m\cup\delta m,$ (2.10) and therefore, up to surface terms, $\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]$ $\to\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]+8\pi^{2}\phi\cup m\cup\delta z-4\pi^{2}\delta\ell\cup m\cup a+8\pi^{3}\mathbb{Z}.$ (2.11) Under the $\mathbb{R}^{(0)}$ gauge transformation (2.9), up to surface terms, $\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]$ $\to\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]$ $\qquad{}+4\pi\left(\phi\cup\delta\lambda-2\pi\ell\cup\lambda\right)\cup\delta z+2\pi\delta\ell\cup\left[\lambda\cup\delta a+2\pi\left(\lambda\cup z+z\cup\lambda+\lambda\cup_{1}\delta z\right)\right].$ (2.12) provided that we set the transformation laws of the auxiliary fields as, under the $\mathbb{Z}^{(1)}$ gauge transformation,
$\tilde{a}\to\tilde{a}+\mathrm{e}^{2}\phi\cup m,\qquad\chi\to\chi-\frac{\mathrm{e}^{2}}{2}m\cup a,$ (2.13)
and, under the $\mathbb{R}^{(0)}$ gauge transformation,
$\tilde{a}\to\tilde{a}+\frac{\mathrm{e}^{2}}{2\pi}\left(\phi\cup\delta\lambda-2\pi\ell\cup\lambda\right),\qquad\chi\to\chi+\frac{\mathrm{e}^{2}}{4\pi}\left[\lambda\cup\delta a+2\pi\left(\lambda\cup z+z\cup\lambda+\lambda\cup_{1}\delta z\right)\right].$ (2.14)
2.2 ’t Hooft line operator as a defect boundary
We introduce the ’t Hooft line operator in the present system as follows. We couple a background 2-cochain gauge field $B\in\mathbb{Z}$ to the electric 2-cochain current, $\star f$. For this, we change the kinetic term of the gauge field in Eq. (2.1) as $f\cup\star f\to(f-2\pi B)\cup\star(f-2\pi B)$ and correspondingly define
$S^{B}$ $:=\sum_{h\in\Gamma}\biggl{\{}\frac{1}{g_{0}^{2}}\left(f-2\pi B\right)\cup\star\left(f-2\pi B\right)+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\partial\phi\cup\star\partial\phi+i\delta\ell\cup\chi$
$\qquad\qquad{}-i\frac{\mathrm{e}^{2}}{8\pi^{2}}\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]\biggr{\}}.$ (2.15)
The Poincaré dual of $B$ defines a 2-dimensional surface (i.e., defect) $\mathcal{R}$. If $\delta B\neq 0$, the surface $\mathcal{R}$ has the boundary $\gamma$, $\partial\mathcal{R}=\gamma$, along which $\delta B\neq 0$. Then the gauge coupling (2.15) defines a ’t Hooft line operator along $\gamma$. Having these facts in mind, we set
$B=-\mathrm{q}\delta_{2}[\mathcal{R}],$ (2.16)
where $\mathrm{q}$ is an integer (magnetic charge) and $\delta_{2}[\mathcal{R}]$ is the 2-cochain delta function whose support is $\mathcal{R}$. This implies that
$\delta B=-\mathrm{q}\delta_{3}[\gamma],$ (2.17)
where $\delta_{3}[\gamma]$ is the 3-cochain delta function along $\gamma$.
In the present axion QED, $\star f$ is not conserved under the equation of motion, $\delta\star f\neq 0$, and thus the partition function is not invariant under the 1-form gauge transformation on $B$, $B\to B+\delta\omega$. Correspondingly, the Poincarè dual of $B$, the surface $\mathcal{R}$ is not topological, i.e., the partition function with the Boltzmann weight $e^{-S^{B}}$ depends on a precise shape of $\mathcal{R}$. Nevertheless, we may still introduce such a defect and define a (non-genuine) ’t Hooft line operator. Note that, since the combination $(f-2\pi B)\cup\star(f-2\pi B)$ is manifestly invariant under the gauge transformations in Eqs. (2.7)–(2.9), the above method defines an object invariant under those gauge transformations. [^5] The continuum counterpart of Eq. (2.1) possesses an electric $\mathbb{Z}_{\mathrm{e}^{2}}$ 0-form symmetry Hidaka:2020iaz ; Hidaka:2020izy and one can consider a genuine ’t Hooft line operator with a fractional magnetic charge as the boundary of the corresponding symmetry operator. Such a fractionally-charged magnetic monopole worldline cannot be considered in the present lattice formulation with the integer $z$. One would have to consider the “excision method” Abe:2023uan to incorporate such a fractionally-charged ’t Hooft line; the excision method has not completely been developed for 4D gauge theory.
If we change the variable $z$ in $S^{B}$ (2.15) as $z\to z+B$, we may remove the background gauge field $B$ in the gauge kinetic term. Instead, the other parts in the action (2.15) produce new factors. Under the shift $z\to z+B$, we find that the Boltzmann weight changes as
$e^{-S^{B}}\to e^{-S}T_{\mathrm{q}}(\gamma),$ (2.18)
where the ’t Hooft line operator $T_{\mathrm{q}}(\gamma)$ is given by
[…]
2.3 Axial $U(1)$ Ward–Takahashi identity
Next, we define the expectation value in the sub-sector by [^11] We assume that, divided by the gauge volumes, the integration measure is defined by $\int\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\tilde{a}(l)\right]\prod_{s\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\phi(s)\right]\prod_{l\in\Gamma}\left[\sum_{\ell(l)=-\infty}^{\infty}\right]\prod_{p\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}\chi(p)\right].$ (2.20)
$\left\langle\dotsb\right\rangle_{\mathrm{s}}^{B}:=\int\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]\,e^{-S^{B}}\dotsb.$ (2.21)
Note that the gauge fields ($a$ and $z$) are not yet integrated in this expression. We then consider a constant shift of the integration variable $\phi$ in a 4-dimensional region $\mathcal{V}_{4}\subset\Gamma$,
$\phi\to\begin{cases}\phi+\alpha&\text{for $x\in\mathcal{V}_{4}$},\\ \phi&\text{otherwise}.\\ \end{cases}$ (2.22)
We assume that $\mathcal{V}_{4}$ contains the 2-surface $\mathcal{R}$ in Eq. (2.16) as Fig. 1.
$\mathcal{M}_{3}^{\prime}$$\mathcal{M}_{3}$$\mathcal{V}_{4}$$\gamma$$\mathcal{R}$
Figure 1:
[…]
$S^{B}$ $\to S^{B}-\frac{i}{2}\alpha\left\{\sum_{c\in\partial\mathcal{V}_{4}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]-\mathrm{e}^{2}\sum_{h\in\mathcal{V}_{4}}P_{2}(z)\right\},$ (2.23)
where $j_{5}$ is the axial $U(1)$ 1-cochain current,
$j_{5}:=-2i\mu^{2}\partial\phi.$ (2.24)
Since the integration measure (2.20) has an invariant meaning under Eq. (2.21), [^12] That is, we may start with the measure $\int_{-\pi+\alpha}^{\pi+\alpha}\mathrm{d}\phi(s)$ for $s\in\mathcal{V}_{4}$ instead of Eq. (2.20) because of the $\mathbb{Z}^{(0)}$ gauge invariance. we have the identity,
$\left\langle\dotsb\right\rangle_{\mathrm{s}}^{B}=\left\langle\exp\left(\frac{i}{2}\alpha\left\{\sum_{c\in\partial\mathcal{V}_{4}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]-\mathrm{e}^{2}\sum_{h\in\mathcal{V}_{4}}P_{2}(z)\right\}\right)\dotsb\right\rangle_{\mathrm{s}}^{B},\qquad\delta z=0.$ (2.25)
[…]
2.4 Final steps
We now assume that in Eq. (2.25) the boundary of $\mathcal{V}_{4}$ consists of two 3-surfaces,
$\partial\mathcal{V}_{4}=\mathcal{M}_{3}^{\prime}\cup(-\mathcal{M}_{3}).$ (2.26)
[…]
$\exp\left\{\frac{i}{2}\alpha\sum_{c\in\mathcal{M}_{3}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]\right\}.$ (2.27)
This is possible, because it can be seen that $\star j_{5}$ on $\mathcal{M}_{3}$ is invariant under Eq. (2.22). Thus, we have
$\left\langle\exp\left\{\frac{i}{2}\alpha\sum_{c\in\mathcal{M}_{3}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})(z)\right]\right\}\right\rangle_{\mathrm{s}}^{B}.$
$=\exp\left[-\frac{i}{2}\alpha\mathrm{e}^{2}\sum_{h\in\mathcal{V}_{4}}P_{2}(z)\right]\left\langle\exp\left\{\frac{i}{2}\alpha\sum_{c\in\mathcal{M}_{3}^{\prime}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]\right\}\right\rangle_{\mathrm{s}}^{B},\qquad\delta z=0.$ (2.28)
Finally, we set
$\alpha=\frac{2\pi p}{N},$ (2.29)
where $p$ and $N$ are coprime integers and define the symmetry operator of the axial $U(1)$ non-invertible symmetry by Honda:2024yte ; Honda:2024sdz
$U_{2\pi p/N}(\mathcal{M}_{3}):=\exp\left\{\frac{i\pi p}{N}\sum_{c\in\mathcal{M}_{3}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]\right\}\mathcal{Z}_{\mathcal{M}_{3}}[z],$ (2.30)
where $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ is the partition function of a lattice 3D TQFT. In Refs. Honda:2024yte ; Honda:2024sdz , an explicit form of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$ on the lattice is given by a level $N$ BF theory. The basic property of $\mathcal{Z}_{\mathcal{M}_{3}}[z]$, for the configuration (2.26) is, for $\delta z=0$ Honda:2024yte ,
$\mathcal{Z}_{\mathcal{M}_{3}}[z]=\exp\left[\frac{i\pi p}{N}\mathrm{e}^{2}\sum_{h\in\mathcal{V}_{4}}P_{2}(z)\right]\mathcal{Z}_{\mathcal{M}_{3}^{\prime}}[z].$ (2.31)
Combining Eqs. (2.28) and (2.31), we thus have for the symmetry operator (2.30),
$\left\langle U_{2\pi p/N}(\mathcal{M}_{3})\right\rangle_{\mathrm{s}}^{B}=\left\langle U_{2\pi p/N}(\mathcal{M}_{3}^{\prime})\right\rangle_{\mathrm{s}}^{B}.$ (2.32)
We may integrate both sides of Eq. (2.32) by $a$ and $z$ to yield the full expectation value. Then, we may shift the integration variable $z$ by $z\to z+B$ to make the presence of the ’t Hooft line explicit. Recalling the relation (2.18), we have
$\left\langle U_{2\pi p/N}(\mathcal{M}_{3})T_{\mathrm{q}}(\gamma)\right\rangle=\left\langle U_{2\pi p/N}(\mathcal{M}_{3}^{\prime})T_{\mathrm{q}}(\gamma)\right\rangle,$ (2.33)
where the full expectation value is defined by [^14] We understand $\int\mathrm{D}[a]\mathrm{D}[z]:=\prod_{l\in\Gamma}\left[\int_{-\pi}^{\pi}\mathrm{d}a(l)\right]\prod_{p\in\Gamma}\left[\sum_{z(p)=-\infty}^{\infty}\right].$ (2.34)
$\left\langle\dotsb\right\rangle:=\int\mathrm{D}[a]\mathrm{D}[z]\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]\,e^{-S}\dotsb.$ (2.35)
We note that the shift $z\to z+B$ does not affect the symmetry operators in Eq. (2.32), because the defect $B$ does not have the support on the 3D boundaries $\mathcal{M}_{3}$ and $\mathcal{M}_{3}^{\prime}$ on which the symmetry operators are defined. Equation (2.33) and Fig. 1 show that the sweep of the non-invertible symmetry operator over the ’t Hooft line operator leaves no effect.
Let us summarize the essential elements in our argument. First, we assumed that the non-genuine but gauge invariant ’t Hooft line operator with an integer magnetic charge can be realized by the defect $B$ (2.16) introduced in the gauge field kinetic term. Next, we assume that the 4D topological part in the axial anomaly $P_{2}(z)$ with the fractional coefficient can be represented by a 3D TQFT as in Eq. (2.31); this enables one to define the symmetry operator as Eq. (2.30). Then, we have the relation (2.33). As far as these assumptions are fulfilled, the above conclusion follows almost independently of the details of the lattice action (i.e., possible modifications of $j_{5}$, $(\mathrm{CS})$, $P_{2}(z)$, and the gauge transformation laws).
2.5 Axion string
For the axion string, the argument is quite parallel. We introduce a 2D defect $\sigma$ by the substitution, $\partial\phi\to\partial\phi-2\pi C$, in the scalar kinetic term as
$S^{C}$ $:=\sum_{h\in\Gamma}\biggl{\{}\frac{1}{g_{0}^{2}}f\cup\star f+i\tilde{a}\cup\delta z+\frac{\mu^{2}}{2}\left(\partial\phi-2\pi C\right)\cup\star\left(\partial\phi-2\pi C\right)+i\delta\ell\cup\chi$
$\qquad\qquad{}-i\frac{\mathrm{e}^{2}}{8\pi^{2}}\left[\partial\phi\cup(\mathrm{CS})-4\pi^{2}\phi\cup P_{2}(z)\right]\biggr{\}},$ (2.36)
where, assuming $\sigma=\partial\mathcal{V}_{3}$, we set
$C:=-\mathrm{q}\delta_{1}[\mathcal{V}_{3}],\qquad\delta C=-\mathrm{q}\delta_{2}[\sigma],$ (2.37)
where $\delta_{1}[\mathcal{V}_{3}]$ is the 1-cochain delta function whose support is $\mathcal{V}_{3}$. Under the change of variable, $\ell\to\ell+C$, we have
$e^{-S^{C}}\to e^{-S}S_{\mathrm{q}}(\sigma),$ (2.38)
where the axion string operator $S_{\mathrm{q}}(\sigma)$ is given by
$S_{\mathrm{q}}(\sigma)=\exp\left\{-i\sum_{h\in\Gamma}\delta C\cup\left[\chi+\frac{\mathrm{e}^{2}}{4\pi}a\cup a\right]+i\sum_{h\in\Gamma}\frac{\mathrm{e}^{2}}{4\pi}C\cup\left(f\cup a+2\pi a\cup z\right)\right\},$ (2.39)
where we have used $\delta z=0$, which is imposed under the integration over $\tilde{a}$. This expression with Eq. (2.37) precisely coincides with that of the axion string operator in Ref. Honda:2024sdz .
We then introduce the expectation value,
$\left\langle\dotsb\right\rangle_{\mathrm{s}}^{C}:=\int\mathrm{D}[\tilde{a}]\mathrm{D}[\phi]\mathrm{D}[\ell]\mathrm{D}[\chi]\,e^{-S^{C}}\dotsb.$ (2.40)
Then, under the shift (2.22),
$\left\langle\dotsb\right\rangle_{\mathrm{s}}^{C}=\left\langle\exp\left(\frac{i}{2}\alpha\left\{\sum_{c\in\partial\mathcal{V}_{4}}\left[\star j_{5}-\frac{\mathrm{e}^{2}}{4\pi^{2}}(\mathrm{CS})\right]-\mathrm{e}^{2}\sum_{h\in\mathcal{V}_{4}}P_{2}(z)\right\}\right)\dotsb\right\rangle_{\mathrm{s}}^{C},$ (2.41)
where we have noted that although the axial $U(1)$ 1-cochain current is modified as $j_{5}\to-2i\mu^{2}\left(\partial\phi-2\pi C\phi\right)$ in the presence of the background $C$, since we assume that the support of $C$, $\mathcal{V}_{3}$, does not touch upon the boundary of the 3D region $\partial\mathcal{V}_{4}$, we can neglect this modification in Eq. (2.41). Then, repeating the same argument as that for the ’t Hooft line operator, we have
$\left\langle U_{2\pi p/N}(\mathcal{M}_{3})\right\rangle_{\mathrm{s}}^{C}=\left\langle U_{2\pi p/N}(\mathcal{M}_{3}^{\prime})\right\rangle_{\mathrm{s}}^{C}.$ (2.42)
and, in the full expectation value (2.35), after the shift of the integration variable $\ell\to\ell+C$, we have
$\left\langle U_{2\pi p/N}(\mathcal{M}_{3})S_{\mathrm{q}}(\sigma)\right\rangle=\left\langle U_{2\pi p/N}(\mathcal{M}_{3}^{\prime})S_{\mathrm{q}}(\sigma)\right\rangle.$ (2.43)
Again, the sweep of the symmetry operator over the axion string does not leave any effect.
Acknowledgment
This work was partially supported by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research Grant Number JP23K03418 (H.S.).